OpenStudy (anonymous):

10v^7+28v^6-6v^5 factor

OpenStudy (anonymous):

10v^(7)+28v^(6)-6v^(5) Factor out the GCF of 2v^(5) from the expression 10v^(7). 2v^(5)(5v^(2))+28v^(6)-6v^(5) Factor out the GCF of 2v^(5) from the expression 28v^(6). 2v^(5)(5v^(2))+2v^(5)(14v)-6v^(5) Factor out the GCF of 2v^(5) from the expression -6v^(5). 2v^(5)(5v^(2))+2v^(5)(14v)+2v^(5)(-3) Factor out the GCF of 2v^(5) from 10v^(7)+28v^(6)-6v^(5). 2v^(5)(5v^(2)+14v-3) For a polynomial of the form ax^(2)+bx+c, find two factors of a*c (-15) that add up to b (14).In this problem 3*-(1)/(5)=-(3)/(5) (which is (c)/(a)) and 3-(1)/(5)=(14)/(5) (which is ((b)/(a)) , so insert 3 as the right hand term of one factor and -(1)/(5) as the right-hand term of the other factor. 2v^(5)(v+3)(v-(1)/(5)) Remove the fraction by multiplying the first term of the factor by the denominator of the second term. 2v^(5)(v+3)(5v-1)

OpenStudy (anonymous):

do you fully understand the concept now?

OpenStudy (anonymous):

yes